# An Improved Two-Step Attack on Kyber

### v2

@@ Abstract.
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by the national institute of standards and technology (NIST), CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer energy traces and less time. In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of energy traces. The CPA is enhanced by utilizing both the Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover s based on the results of CPA. The success rate is largely built up by constructing a trial-and-error method. We implement the proposed attack for the reference implementation of Kyber512 (4 128-value groups of s) on ARM Cortex-M4 and successfully recover a 128-value group of s in about $9$ minutes using a $16$-core machine. Additionally, in that case, we only cost at most $60$ CPA guess values for a group and 15 power traces for a guess.
@@
Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 2024 ACM/IEEE International Conference on Computer-Aided Design; October 27–31, 2024; NEW JERSEY, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
WITH the fast development of quantum computing, modern public-key cryptographic algorithms are encountering significant threats. Recognizing this problem, the national institute of standards and technology (NIST) started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms. By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation (alagic2022status, ). CRYSTALS-Kyber is that only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024 (Announcing2024, ). It should be marked that the mathematical security of Kyber has been widely recognized by the cryptography community. However, for the upcoming deployment phase, it becomes urgent to emphasize its physical security (nagata2021physical, ; zhao2023side, ; ravi2024side, ).
Kocher et al. first introduced side-channel attacks (SCAs) in 1996 by leveraging the data dependency on power consumption of cryptographic devices (kocher1996timing, ). Generally, in the absence of any protective measures, devices running the cryptographic algorithms with long-term keys are susceptible to SCAs, especially to the correlation power analysis (CPA).
@@
Recent years, several CPA attacks on Kyber (ravi2020generic, ; ueno2022curse, ; shen2023find, ; yang2023chosen, ; kuo2023lattice, ) have been proposed by researchers. In 2020, Ravi et al (ravi2020generic, ). used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time. Then, Ueno et al (ueno2022curse, ). proposed the side channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key. The work in (shen2023find, ) deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber512 at the cost of about 1619 power traces, but there are still 40% error coefficients. Contrast to previous works, in (yang2023chosen, ), Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber512, successfully recovering two coefficients of secret key using $20$ or more power traces. Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber512. So, they at least need 5120 power traces for a key. In (kuo2023lattice, ), Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of M-LWE problem. The correctness was verified by using computer simulations. In this paper, we take the idea of (kuo2023lattice, ) as the starting point and develop an improved two-step attack method on Kyber to further validate it in practical and improve the efficiency.
Figure 1. The flowchart of the two-step attack.
@@
The proposed two-step attack is summarized in Fig. 1. In the first step, we modify the leakage model to better align our CPA with the arithmetic logic unit. Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks. In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key, and optimize the parameters of the algorithm. By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice.
2. Preliminaries
2.1. NTT in CRYSTALS-Kyber
@@
Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512^{\text{th}}$. Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations. The even coefficients in NTT domian are described as follows:
(1) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$
@@
where $br$($i$) represents a 7-bit bit-reversal of $i$ and $\mathbf{M}$ is a $128\times 128$ matrix with reduced integers. Similarly, INTT can be represented in the same way and the inverse matrix is denoted as $N$.
Listing 1: Base multiplication in Kyber.
1void basemul(int16_t r[2],
2 int16_t zeta,
3 const int16_t s[2],
4 const int16_t u[2])
5{
6 r[0] = fqmul(s[1], u[1]);
7 r[0] = fqmul(r[0], zeta);
8 r[0] += fqmul(s[0], u[0]);
9 r[1] = fqmul(s[0], u[1]);
10 r[1] += fqmul(s[1], u[0]);
11}
2.2. Correlation-Based Attack Techniques
@@
The conventional CPA attack is a non-invasive side-channel attack that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys. It identifies the correct key by finding the highest correlation between measured and hypothetical values. Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
@@
In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together (abdi2007kendall, ). Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement. The Kendall’s tau calculation is computed as follows:
(2) $\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$
@@
3. Proposed Improved Two-Step Attack
@@
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We will detail the basis of the attack analysis and our proposed three improvements in the following. We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device. In addition, they can also capture the power traces of the device.
@@
Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber (shen2023find, ; yang2023chosen, ). One critical side-channel leakage occurs during the point-wise multiplication of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. In the clean implementation of pqm4 (kannwischer2019pqm4, ), polynomial multiplication employs 256 $basemuls$ in Kyber512. Moreover, each calculation of $basemul$ is independent from the others. From Listing 1, it is clear that the result depends on two secret key coefficients, $s_{0}$ and $s_{1}$. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$. $r_{0}$ can be selected as the intermediate value.
3.1. Modified Leakage Model for CPA
@@
The Hamming weight (HW) model counts the number of ”1” in intermediate values and is a common linear power model. ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values. This modification improves the accuracy of the HW model and reduces the number of power traces (denoted as $D$) needed for CPA. As shown inside the blue dotted box in Fig. 2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
3.2. Optional Kendall’s tau for CPA
@@
When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys. To address this, we set a PCC threshold (e.g., 0.9) and extract candidate key coefficients with correlations exceeding this threshold. We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces. The candidate with the highest Kendall’s tau value is selected as the correct key coefficient, which helps to reduce false positives. This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
@@
Figure 2. (a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively. (c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
3.3. Trial-and-Error Lattice Attack after CPA
@@
As introduced in (kuo2023lattice, ), suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered. Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ (kuo2023lattice, ), where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
@@
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We use Kannan’s embedding technique to solve this problem (kannan1987minkowski, ), where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
@@
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128.
But it is worth noting that the norm of vector $\mathbf{w}$ must be less than the norm of the shortest vector estimated by the Gaussian heuristic for the unique shortest vector problem to be solved. For Kyber512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients. For Kyber768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
@@
We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively.
One of the benefits of using the proposed method is that in order to recover a set of key coefficients, we can use a minimum of $38/39$ CPA instead of $128$ CPA. Certainly, the more CPA are carried out, the higher the success rate of this method will be, but its efficiency will also slow down. The one with the smallest correlation among the $38/39$ key coefficients is replaced by other key coefficients with higher correlation than it in a trial-and-error method. It is worth noting that it is necessary to set the appropriate parameters of lattice attack; otherwise, it may take too long in a trial and lead to the overall recovery inefficient. For Kyber512, Kyber768, and Kyber1024, 4, 6, and 8 iterations are needed to recover the full secret key, respectively. Since even-indexed coefficients do not affect odd-indexed coefficients and coefficients within each group do not affect each other, the above method can be parallelized to compute the corresponding coefficients concurrently.
Listing 2: Trail-and-Error Lattice Attack.
1index = [1, 2, ..., 37, 38]
2other_index = [39, 40, ..., n-2, n-1]
3guess_sk = [sk_bu0, ..., sk_bu38]
4other_sk = [sk_bu39, ..., sk_bu(n-1)]
5ken = [k_bu0, ..., k_bu38]
6other_ken = [k_bu39, ..., k_bu(n-1)]
7
8result = Lattice_attack(index, guess_sk)
9result = (M @ result)
10if guess_sk == [result[j] for j in index]:
11 return result
12else:
13 for i in count:
14 updata_index(index)
15 updata_sk(guess_sk)
16 result = Lattice_attack(index, guess_sk)
17 result = (M @ result)
18 result_target = [result[j] for j in index]
19 if guess_sk == result_target:
20 return result
21 break
22
23def Lattice_attack(a, b):
24def updata_index():
25def updata_sk():
4. Experiments

### v3

# An Improved Two-Step Attack on Lattice-based Cryptography: A Case Study of Kyber

@@ Abstract
After three rounds of post-quantum cryptography (PQC) strict evaluations conducted by NIST, CRYSTALS-Kyber was successfully selected in July 2022 and standardized in August 2024. It becomes urgent to further evaluate Kyber’s physical security for the upcoming deployment phase. In this brief, we present an improved two-step attack on Kyber to quickly recover the full secret key, $\mathbf{s}$, by using much fewer power traces and less time. In the first step, we use the correlation power analysis (CPA) to obtain a portion of guess values of $\mathbf{s}$ with a small number of power traces. The CPA is enhanced by utilizing both Pearson and Kendall’s rank correlation coefficients and modifying the leakage model to improve the accuracy. In the second step, we adopt the lattice attack to recover $\mathbf{s}$ based on the results of CPA. The success rate is largely built up by constructing a trial-and-error method. We deploy the reference implementations of Kyber-512, -768, and -1024 on an ARM Cortex-M4 target board and successfully recover $\mathbf{s}$ in approximately $9\sim 10$ minutes with at most 15 power traces, using a Xeon Gold 6342-equipped machine for the attack.
Index Terms:
@@
Lattice-based cryptography, CRYSTALS-Kyber, Side-channel attack, Power analysis.
I Introduction
@@
WITH the fast development of quantum computing, modern cryptographic algorithms are encountering significant threats. Recognizing this problem, NIST started the post-quantum cryptography (PQC) standardization process in 2016 with the aim of standardizing quantum-resistant cryptographic algorithms. By July 2022, NIST released the post-quantum cryptographic standard candidates including three signatures and one key encapsulation mechanism (KEM) algorithm after the third round evaluation [1]. CRYSTALS-Kyber is the only KEM, which has been successfully standardized as module-lattice-based KEM (ML-KEM) noted as FIPS 203 from August 2024. The mathematical security of Kyber is widely recognized, but its physical security demands urgent attention in the deployment phase [2, 3].
@@
In recent years, several side-channel attacks (SCAs) on Kyber [4, 5, 6, 7, 8] have been proposed by researchers. In 2020, Ravi et al. [4] used plaintext-checking (PC) oracle-based SCAs to recover the secret key of Kyber-512 with 2560 power traces according to the vulnerabilities in Fujisaki-Okamoto (FO) transform for the first time. Then for correlation power analysis (CPA), Ueno et al. [5] proposed the side-channel leakage in re-encryption as a PC oracle, utilizing 3072 power traces to recover the key. The work in [6] deals with imperfect PC oracle SCAs by using the CPA attack twice to recover 60% secret key of Kyber-512 at the cost of about 1619 power traces, but there are still 40% error coefficients. Contrast to previous works, in [7], Yang et al. tried to carry out a random ciphertext CPA attack on the reference implementation of Kyber-512, successfully recovering two coefficients of secret key using $20$ or more power traces. Nevertheless, they had to repeat the attack 256 times to recover the entire secret key of Kyber-512. So, they at least need 5120 power traces for a key. In [8], Yen-Ting Kuo and Atsushi Takayasu presented a two-step attack on Kyber by integrating the random ciphertext CPA attack with the lattice attack. They adopted the CPA attack to recover part of the coefficients of a key and then directly solved the rest coefficients by using a low-dimension equation of the M-LWE problem. The correctness was verified by using computer simulations. In this brief, we take the idea of [8] as the starting point and develop an improved two-step attack method on Kyber to further validate it in practice and improve its efficiency.
@@
The proposed two-step attack is summarized in Fig. 1. In the first step, we modify the leakage model to better align our CPA with the arithmetic logic unit. Additionally, we introduce an optional correlation coefficient, the Kendall rank correlation, to address the nonlinearity and robustness issues of the original CPA in real-world attacks. In the second step, we construct a trial-and-error lattice attack algorithm to recover the entire secret key and optimize the algorithm’s parameters. By employing this two-step attack, we achieve significant improvements in both efficiency and accuracy in practice. In addition, the attack methods we proposed for the Point-Wise Multiplication (PWM) stage apply to several other lattice-based PQC. Dilithium employs the single-point PWM and is more vulnerable to similar attacks. The PWM of Falcon works with floating-point numbers, but significant side-channel leakage remains due to secret key and message multiplication [9].
II Preliminaries
II-A Kyber and Number Theoretic Transform in Kyber
Kyber consists of three stages: key generation, encryption, and decryption. In the key generation stage, the public key $pk$ is constructed as $\mathbf{A}\mathbf{s}+\mathbf{e}$. During the encryption stage, the message $m$ is encrypted into ciphertext $c$. In the decryption stage, the receiver utilizes $\mathbf{s}$ to perform corresponding operations to recover the message $m$ from $c$. The main operations of decryption are as follows:
$m\leftarrow Compress_{q}(v-INTT(\boldsymbol{\hat{s}}^{T}\circ NTT(\boldsymbol{u})),1).$ (1)
@@ II-A Kyber and Number Theoretic Transform in Kyber
Computing polynomial multiplication in NTT domain can reduce the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N\log N)$, thus significantly improving the speed of the entire encryption and decryption process in Kyber. Note that for Kyber, here is only $256^{\text{th}}$ primitive root of unity $\xi$ instead of $512^{\text{th}}$. Therefore, 256 coefficients are divided into two groups, namely the odd and even groups, for NTT calculations. The even coefficients in NTT domain are described as follows:
$\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\xi^{(2br(i)+1)j}\Rightarrow\ [\hat{f}_{2i}]^{\top}=\mathbf{M}[f_{2i}]^{\top},0\leq i<128,$ (2)
@@
where $br$($i$) represents bit-reversal of $i$ and $\mathbf{M}$ is an $128\times 128$ matrix with reduced integers. Similarly, the inverse matrix of INTT can be represented as $\mathbf{N}$ in the same way.
Figure 1: The architecture of the proposed two-step attack.
II-B Correlation-Based Attack Techniques
@@
The conventional CPA attack is a non-invasive SCA that uses the Person correlation between power consumption models and actual power traces to recover cryptographic keys. It identifies the correct key by finding the highest correlation between measured and hypothetical values. Pearson correlation coefficient (PCC) is efficient and well-suited for detecting linear relationships between the hypothetical and actual values.
@@
In contrast, Kendall’s tau is a non-parametric measure of association that quantifies the degree of concordance between the rankings of two variables, indicating whether their values are consistently ranked together [10]. Kendall’s tau $\tau$ ranges from $-1$ to $1$, with $\tau=1$ indicating perfect agreement, $\tau=-1$ indicating perfect disagreement, and $\tau=0$ indicating no agreement. The Kendall’s tau calculation is computed as follows:
$\tau_{b}=\frac{c-d}{\sqrt{({c+d+t_{x}})({c+d+t_{y}})}},$ (3)
@@
III Proposed Improved Two-Step Attack
@@
It should be remarked that the proposed improved two-step attack is illustrated in Fig. 1. We assume that an attacker has access to a device which is running Kyber decryption and can only observe ciphertexts entered into the device. In addition, they can also capture the power traces of the device.
@@
Several test vector leakage assessment (TVLA) results from previous works have indicated the potential leakage points during the decryption process in Kyber [6, 7]. One critical side-channel leakage occurs during the PWM of $\hat{\mathbf{s}}\circ\hat{\mathbf{u}}$. In the clean implementation of pqm4 [11], polynomial multiplication employs 256 $basemuls$ in Kyber-512. Moreover, each calculation of $basemul$ is independent from the others. From Listing 1, it is clear that the result depends on two secret key coefficients, $s_{0}$ and $s_{1}$. All partial secret key coefficients range from $\left(0,3328\right)$ for Kyber. In a practical attack, suppose that the attacker aims to recover the coefficient $s_{0}$. $r_{1}$ can be selected as the intermediate value.
III-A Modified Leakage Model for CPA
@@
The Hamming weight (HW) model counts the number of “1” in intermediate values and is a common linear power model. ARM processors’ ALUs are optimized for two’s complement operations, so we adjust the HW (called modified Hamming weight, MHW) to account for this by using the HW of the 16-bit complement of the intermediate values. This modification improves the accuracy of the HW model and reduces the number of power traces (denoted as $D$) needed for CPA. As shown inside the blue dotted box in Fig. 2 (a) and (b), the correct key coefficient is obscured using the original model (a) but can be effectively identified with just 15 power traces using the modified model (b).
III-B Optional Kendall’s tau for CPA
@@
When the number of power traces is limited, traditional CPA may face challenges due to high PCC values across guessed keys. To address this, we set the ||||PCC|||| threshold (e.g., 0.9) based on repeated experimental analysis, with 0.9 performing well in most $basemul$ calls. We then map these candidates to a hypothetical power consumption matrix and compute Kendall’s tau between this matrix and the actual power traces. The candidate with the highest Kendall’s tau value is selected as the correct key coefficient, which helps to reduce false positives. This approach combines PCC for initial filtering with Kendall’s tau for more precise selection, despite Kendall’s tau being more computationally demanding.
@@
Figure 2: (a) and (b) present the PCC values for all guessed keys in a $basemul$ call when $D=15$, for the original and modified models, respectively. (c) shows the ||||PCC|||| values for all guessed keys, while the blue curve in (d) illustrates the occurrence of false positives.
III-C Trial-and-Error Lattice Attack after CPA
@@
As introduced in [8], suppose that an attacker has successfully recovered $sr$ coefficients out of a group of 128 secret key coefficients $\hat{\mathbf{s}}_{i}$, while the remaining $128-sr$ coefficients are unsuccessfully recovered. Let $I_{a}=(a_{0},\ldots,a_{sr-1})$ denote the recovered coefficients indices, and $I_{b}=(b_{0},\ldots,b_{127-sr})$ represent the indices of coefficients that have failed to be recovered, i.e., the unknown coefficients. INTT($\hat{\mathbf{s}}_{i}$) = $\mathbf{N}\hat{\mathbf{s}}_{i}$ = $\mathbf{s}_{i}$ mod $q$ can be rewritten as $\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$ + $\mathbf{N}_{B}\hat{\mathbf{s}}_{iB}$ = $\mathbf{s}_{i}$ mod $q$ [8], where matrix $\mathbf{N}_{A}=[\mathbf{n}_{a_{0}},\ldots,\mathbf{n}_{{a}_{sr-1}}]$ consists of columns in matrix $\mathbf{N}$ corresponding to the indices $I_{c}$, while $\hat{\mathbf{s}}_{iA}=[\hat{\mathbf{s}}_{a_{0}},\ldots,\hat{\mathbf{s}}_{{a}_{sr-1}}]^{\top}$ represents the vector composed of successfully recovered coefficients. Similarly, $\mathbf{N}_{B}$ and $\hat{\mathbf{s}}_{iB}$ are obtained in the same manner.
@@
In the above formulas, both $\mathbf{N}_{A}$ and $\hat{\mathbf{s}}_{iA}$ are known. Let $\mathbf{t}=\mathbf{N}_{A}\hat{\mathbf{s}}_{iA}$, $\mathbf{A}=-\mathbf{N}_{B}$, and $\mathbf{s}^{\prime}=\hat{\mathbf{s}}_{iB}$. Then, we obtain $\mathbf{t}=\mathbf{A}\cdot\mathbf{s}^{\prime}+\mathbf{s}_{i}\mod q$. This conveniently forms a low-dimension LWE problem, which is simpler compared to the original problem in Kyber because the rank of $\mathbf{A}$ is smaller. We use Kannan’s embedding technique to solve this problem [12], where the matrix $\mathbf{B}_{kan}$ is summarized as follows:
@@
$\left[\mathbf{I}_{sr}\mid\mathbf{A}^{\prime}\right]$ is a reduced row echelon matrix of $\mathbf{A}$ transpose. A vector $\mathbf{w}=[\mathbf{s}_{i}^{\top}|1]$ of length $129$ is obtained by reduction algorithm, which contains the secret key $\mathbf{s}_{i}^{\top}$ of length 128. For Kyber-512, when $sr\geq 39$, a lattice attack can recover 128 odd/even index coefficients. For Kyber-768/1024, when $sr\geq 38$, the corresponding coefficients can be correctly recovered.
@@
We elaborate on the theoretical foundation of lattice attacks, which can tolerate error coefficients of recovered key and still successfully retrieve all 128 key. Leveraging this method, we propose a more flexible and versatile approach for its application. The proposed trial-and-error lattice attack is shown in Listing 2, where $n$ represents the number of CPA, and $\mathbf{M}$ and $count$ denote the NTT linear matrix form of Kyber and the number of trials, respectively. In our method, we select 60 basemul calls for the CPA attack and use 38/39 of them as inputs for the lattice attack. If the attack fails, the smallest input is replaced with another CPA result. This process continues until the attack succeeds or all 60 results are used, representing a trade-off between computational complexity and success rate. After extensive experimentation, we found that using 60 results provides similar success to using 128. It is also important to note that as the dimension of the lattice problem increases, the computational complexity grows. For Kyber, 38/39 results are sufficient to handle the 128 coefficients of a key group. However, for Dilithium with a group size of 256, both the lattice dimension and the required number of CPA results will need to increase.
IV Experiments